OpenStudy (dobby1):

ANYONE GOOD AT MATH standard form of a line with a slope of 7/8 and passes through points (2, 7) and (–6, 0)

OpenStudy (dobby1):

standard form of a line with a slope of 7/8 and passes through points (2, 7) and (–6, 0)

OpenStudy (dobby1):

@amistre64 @ash2326

OpenStudy (dobby1):

I worked it out to this point but can't figure out what to do next

OpenStudy (mokeira):

y=mx+c take any one point and the slope \[\frac{ 7 }{8 }=\frac{ y-0 }{ x--6 }\] \[\frac{ 7 }{8 }=\frac{ y }{ x+6 }\] cross multiply \[7(x+6)=8y\] simplify and arrange in the form of y=mx+c

OpenStudy (mokeira):

where did you reach in your working...we can try work from where you got stuck

OpenStudy (dobby1):

I found the slope and was stuck

OpenStudy (mokeira):

oh, ok...have you understood what i have done?

OpenStudy (dobby1):

Am I correct in saying that I can just simplify your final equation above for the answer?

OpenStudy (mokeira):

definately,that is what you should do :) i have left you to do the rest lol

OpenStudy (dobby1):

its ok thx so much I just started pre calc and am still learning how to do some stuff I am not usually like this with math but I guess after this past summer I need to get myself back in a groove

OpenStudy (mokeira):

lol...no worries

OpenStudy (dobby1):

thx so much if you ever find you need help witch I doubt but if you ever do mention me and I will see what I can do

OpenStudy (mokeira):

hahaha lol!! i do need help sometimes and i will mention you..soon!! so better get yourself back in a groove lol

OpenStudy (dobby1):

yeah I will

OpenStudy (dobby1):

@Mokeira

OpenStudy (dobby1):

the final product of your equation is not in my answers

OpenStudy (mokeira):

what are the answers?

OpenStudy (dobby1):

–7x + 8y = 42 –7x – 8y = 42 –7x + 8y = –42 8x + 7y = 42

OpenStudy (mokeira):

what did you get when you simplified.... the answer is among the choices

OpenStudy (dobby1):

7x+42=8x

OpenStudy (mokeira):

\[7(x+6)=8y\] \[7x+42=8y\] take 7x to the other side \[42=8y-7x\] which is same as \[-7x+8y=42 \]

OpenStudy (dobby1):

oh ok thx s much for coming back and helping me

OpenStudy (mokeira):

u elcome :)

OpenStudy (mokeira):

welcome*